# SySLLM: Generating Synthesized Policy Summaries for Reinforcement Learning Agents Using Large Language Models

### v1

@@ 7.4 Results
For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences. Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ ($t=-0.6202$, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$). However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
@@ 8 Discussion and Future Work
We introduce SySLLM, leveraging LLMs to generate summaries of reinforcement learning agents. Its effectiveness is demonstrated across five distinct environments, including seven different agents. Expert evaluations show that SySLLM closely aligns with expert-generated summaries, capturing key aspects of agent behavior. User studies reveal that participants rate SySLLM summaries higher than HIGHLIGHTS on subjective metrics and perform equally well in agent identification tasks. These results provide evidence for the potential usefulness of the proposed approach. We now turn to discuss several limitations of the SySLLM method and possible ways to address them, as well as additional directions for future work.
@@
SySLLM relies on the availability of a captioning function, a widely accepted approach in language-integrated methodologies, as evidenced by prior research [13, 26, 20]. In simulated environments, captioning is particularly straightforward, as ground truth states are readily accessible, and numerous RL environments, such as NetHack/MiniHack [32, 22] and text-based games [35, 11], already incorporate language-based features. In more complex or real-world scenarios, robust techniques such as object detection [41], cutting-edge captioning frameworks [36], action recognition systems [21], or advanced vision-language models [2] can be utilized. Alternatively, annotator models [13] or pretrained foundation models [37] offer viable solutions for generating captions. With continued advancements in general-purpose captioning models, we anticipate that readily available, off-the-shelf captioning tools will soon support an even broader range of tasks, further solidifying the practicality of this approach.
Advancements in LLMs, particularly larger context windows, can extend this approach to large-scale and more complex environments, further validating its effectiveness and unlocking its potential to provide deeper insights into agent behavior.
Future work could explore a question-answering (QA) framework, allowing users to query the agent’s policy directly. For example, users could ask why specific actions were taken or predict the agent’s behavior in new scenarios. This interactive system would provide deeper insights into the agent’s decision-making process, offering a flexible and intuitive way to understand its behavior.
Another avenue for future work is generating multimodal summaries, combining natural language explanations with visual representations of key moments, trajectories, and decisions. That is, for each pattern identified by the LLM, the summary would also include demonstrations of the pattern extracted from agent traces. This approach could enhance user understanding by integrating text and visuals to more effectively convey complex policies. Visual elements from the agent’s experience dataset might help mitigate hallucinations by grounding explanations in observable behavior, potentially improving reliability and trustworthiness.
@@ Appendix D Unlock Goal-Directed Agent Full Summary
Appendix E Implementation Details
@@
Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
@@
Appendix F Experts instructions:
@@
Appendix G Scale Used in Task 1
@@
Appendix H Systematic Exploration of the Prompt Design

### v2

@@ 7.4 Results
For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences. Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ (t=0.6202t=-0.6202, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$). However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
@@ 8 Discussion and Future Work
In this paper, we introduced SySLLM, an innovative framework leveraging large language models to synthesize reinforcement learning policies into clear, structured narratives. Our method addresses key shortcomings of traditional extractive-visual methods by moving away from fragmented, scenario-specific snapshots toward holistic, language-based summaries. This shift significantly reduces the effort required for users to interpret complex agent behaviors, facilitating better comprehension of strategic decisions and adaptability. Our evaluations provide strong evidence that SySLLM effectively captures nuanced policy behaviors and aligns closely with expert-generated summaries. The qualitative alignment with human experts highlights the capability of language models to accurately convey intricate behavioral patterns and strategic nuances, reinforcing the effectiveness of a textual-abstractive approach. Furthermore, user preference studies demonstrate clear advantages in interpretability and user engagement over visual-based summarization methods.
The broader implications of SySLLM suggest promising new directions for research and practice. We advocate for increased exploration into abstractive-textual explanations in global explanation in reinforcement learning, encouraging the community to investigate advanced techniques such as fine-tuning, in-context learning, and integration with reasoning frameworks. Embracing this paradigm can lead to significant advancements in how we communicate complex policies, ultimately enhancing trust, transparency, and adoption of reinforcement learning in critical real-world scenarios.
Captioners
@@
Our approach relies on the availability of effective captioning functions, a common requirement in language-integrated methodologies as seen in prior research [35, 46, 45]. In simulated environments, this dependency is relatively straightforward to address, as ground truth states are accessible, and many RL environments—such as NetHack/MiniHack [56, 57] and text-based games [58, 59]—already incorporate language-based features. For more complex or real-world scenarios, however, implementations would require robust object detection [60], advanced captioning frameworks [61], action recognition systems [62], or vision-language models [63]. Alternatively, annotator models [35] or pretrained foundation models [64] could generate captions. The continued advancement of general-purpose captioning models suggests that off-the-shelf solutions will soon support an even broader range of tasks, enhancing the practicality of our approach.
LLMs Limitations
The current performance of SySLLM is bounded by the capabilities and limitations of the underlying language models. Context window constraints restrict the amount of agent trajectory data that can be analyzed at once, potentially limiting the ability to identify long-term patterns or rare behaviors. Additionally, LLMs can occasionally generate plausible-sounding but incorrect explanations (hallucinations), particularly when faced with unusual or complex agent behaviors. As LLM technology evolves, particularly with expanded context windows and improved factuality, we anticipate that these limitations will diminish, enabling more comprehensive and reliable agent behavior summaries.
Building on SySLLM’s strengths while addressing its limitations, we identify several promising directions for future research:
Interactive Question-Answering Framework
A promising avenue for future work is extending SySLLM to support an interactive question-answering (QA) framework. This would enable users to directly query the agent’s behavior, asking targeted questions such as “Why did the agent choose action A over B in this situation?” or “How would the agent respond to condition X?” Such interactivity could transform SySLLM from a static summarization tool into a dynamic interface for policy exploration, offering a more flexible and intuitive means of understanding agent behavior. Prior work on logical querying of agents, such as [65], highlights the feasibility and potential of QA-based explanations in RL. Inspired by these developments, we envision a natural language-based QA interface grounded in narrative trajectories, allowing users to probe specific aspects of agent decision-making in a more accessible way.
Multi-modal Summaries
Integrating natural language explanations with visual representations would significantly enhance the communicative power of SySLLM. For each behavioral pattern identified by the language model, the system could provide corresponding visual evidence extracted from agent traces such as state transitions, action sequences, or critical decision points. This multi-modal approach would ground abstract explanations in concrete examples, potentially mitigating hallucination issues while making complex policies more accessible to users with diverse backgrounds and expertise levels.
Comparative Analysis Framework
Future work could extend SySLLM to automatically compare multiple agents’ behaviors, identifying similarities, differences, and relative strengths. This comparative capability would be particularly valuable for evaluating alternative training approaches, understanding the effects of hyperparameter changes, or analyzing how agent behavior evolves during training. By articulating these comparisons in natural language, SySLLM could provide intuitive insights into the factors that influence agent performance and behavior.
@@ Appendix D Unlock Goal-Directed Agent Full Summary
Appendix E Crafter Summarization Prompts
You are analyzing multiple episode summaries of an agent’s behavior in the Crafter environment. Your goal is to identify consistent behavioral patterns across episodes and develop an objective behavioral profile.
[Cross-Episode Analysis Framework]
Systematically evaluate these key dimensions:
Behavioral Consistency
- Identify actions and strategies that appear consistently across episodes
- Calculate the variance in action distributions between episodes
- Note any evolution or change in behavior across sequential episodes
- Determine if the agent displays consistent preferences or purely situational responses
Achievement Patterns
- Calculate achievement unlock rate and consistency across episodes
- Identify which achievements are most frequently obtained
- Analyze the typical sequence or prerequisites leading to achievements
- Assess whether achievement patterns suggest intentional pursuit or incidental acquisition
Resource Priorities
- Identify primary resources consistently targeted across episodes
- Analyze typical crafting sequences when resources are available
- Evaluate how consistently the agent manages inventory
- Determine if there are clear resource collection preferences
Environmental Interaction Patterns
- How consistently does the agent navigate the environment?
- Identify common responses to specific environmental features
- Analyze patterns in exploration vs. exploitation behavior
- Evaluate adaptation to threats, opportunities, and constraints
Decision-Making Characteristics
- Identify the apparent decision criteria for different action choices
- Analyze how the agent balances short-term vs. long-term needs
- Evaluate how predictable the agent’s responses are to similar situations
- Assess whether actions appear purposeful or random
[Output Instructions]
1. Begin with a "Behavioral Profile" summarizing the agent’s most consistent traits
2. Include a "Statistical Analysis" section with quantitative breakdowns of action patterns
3. Provide a "Decision Pattern Analysis" detailing how the agent makes choices
4. Add an "Achievement Analysis" showing typical patterns in achievement progression
5. Conclude with "Behavioral Consistency Assessment" that evaluates how predictable the agent is
Give the agent a label based on its observed behavior and justify your choice. Your analysis should be based entirely on observable patterns. If the agent shows highly inconsistent behavior across episodes, explicitly detail this with supporting evidence. Focus on describing what the agent does consistently, rather than speculating on why it might do so.
Appendix F Implementation Details
F.1 MiniGrid
@@
Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$
@@
F.2 Crafter
We implemented DreamerV3 for our agent, using a state-of-the-art world model-based reinforcement learning approach. The agent processes $64\times 64\times 3$ RGB observations from the Crafter environment. The world model consists of three key components: an encoder network, a recurrent state-space model (RSSM), and a decoder network. The encoder transforms raw pixel observations into a 1024-dimensional embedding space using a convolutional neural network with a depth of 96 channels.
The RSSM, which forms the core of the agent’s predictive capabilities, utilizes a deterministic state of dimension 4096 and a stochastic state represented as a 32-dimensional random variable, allowing the agent to account for environment stochasticity. For temporal dynamics, we employed a GRU cell with 1024 hidden units. The decoder reconstructs observations using transposed convolutions, enabling the model to learn compact state representations through reconstruction loss.
For policy learning, we used an actor-critic architecture with 5-layer MLPs for both actor and critic, where the actor employs a categorical distribution over the 17 discrete actions available in Crafter. The agent was trained using the ”reinforce” gradient strategy for imagination-based policy optimization, with a $\lambda$-return horizon of 15 steps and a discount factor of 0.997.
Training was conducted for $10^{6}$ environment steps using 8 parallel environments, with a batch size of 32 and sequence length of 64. We employed a model learning rate of $10^{-4}$ and an actor learning rate of $3\times 10^{-5}$, optimized using Adam.
Appendix G Experts instructions:
@@
Appendix H Scale Used in Task 1
@@
Appendix I Systematic Exploration of the Prompt Design